Evaluate the double integral || f(x, y) dA over the region D. JU f(x, y) = 6x + 9y and D = {(x, y)SXS 1, x3 sy s x3 + 1}
Answer #1

IdA ; D = {(x,y)] <ulan² = y = x² +1}; frv,y)= 6u+9y U Sfrm,y) da - If Henry) dyda 23 +1 ✓ J (64+94) dy dre 1 - So y =43 N=0= (3x2 + 4x +421.3 = 3(192 +40134+40 -3012-7607-200) = 3+4+4-0 = 12+9+18 = 39 :: AM - -19- CSScanned with CamScanner